# Influence of microscopic parameters on phase behavior of a cell model with Curie-Weiss interaction

### v1

@@ Abstract
We have investigated the effect of changing two microscopic parameters in a cell model with Curie-Weiss-type interaction on its phase behavior, namely the cell volume and ratio between repulsion and attraction intensities. The results are based on an exact solution previously derived for this model in the grand canonical ensemble. At sufficiently low temperatures, the cell model exhibits multiple first-order phase transitions. Varying the cell volume and the repulsion-to-attraction ratio, we represent a quantitative comparison of the chemical potential and pressure isotherms, as well as the pressure-temperature and temperature-density phase diagrams. The analysis of provided data shows that altering microscopic parameters does not lead to qualitative changes in the overall phase behavior of the cell model.
@@ 1 Introduction
The analysis of model fluids using simplified interaction potentials is essential for understanding the connection between microscopic interactions and macroscopic phase behavior. Fluid phase diagrams can be significantly altered by even small changes in interaction parameters, such as well depth, range, or softness of repulsion as shown by computer modeling and theoretical investigations employing statistical mechanics, integral equations, and perturbation theories [1]. Computer simulation studies demonstrate that decreasing the range of attraction in a system of hard spheres with an attractive Yukawa potential [2, 3] as well as in Lennard-Jones like models [4, 5] leads to the vanishing of stable fluid-fluid transitions and remaining stable fluid-solid transitions. Moreover, reducing the interaction range flatten out the vapor-liquid coexistence curves and shrinks the range of temperatures for a stable liquid phase in attractive hard-core Yukawa fluids, as shown by computer modeling [6, 7, 8], reference hypernetted chain integral equation theory [9], and perturbative density functional theory [10]. The same effect is also observed in square-well fluids when decreasing the range of the potential well [11], in systems with interactions modeled by the approximate non-conformal theory potentials when weakening the attractive softness [12], in Mie fluids when either increasing repulsion exponent [13, 14] or increasing dispersive exponent [14], in fluids with modified Buckingham exponential-6 potential when increasing repulsion intensity [15, 16].
@@ 3 Effect of changing the cell size
In this Section, we study the effect of changing the reduced cell volume, $v^{*}$, on the phase behavior of the cell model. In the framework of the cell model the cell volume is a constant quantity. Moreover in the thermodynamic limit the average number of particles $\langle N\rangle\rightarrow\infty$, the total volume of the system $V\rightarrow\infty$, but $v={\mathrm{const}}$. Thus our investigation means the quantitative comparison of the values that the reduced particle number density, $\bar{n}$, the chemical potential, $\tilde{\mu}$, and the pressure, $\tilde{P}$, can take, in the cases we set $v^{*}=1$, $v^{*}=2$, $v^{*}=5$, $v^{*}=10$, $v^{*}=15$.
@@
Figure 1: Part a: effect of changing the cell size, $v^{*}$, on isotherms of the chemical potential, $\tilde{\mu}$, as a function of the reduced density, $\bar{n}$: black line corresponds to $v^{*}=1$, red line – $v^{*}=2$, blue line – $v^{*}=5$, green line – $v^{*}=10$, purple line – $v^{*}=15$. Part b: isotherm of the pressure, $\tilde{P}$, as a function of the reduced density, $\bar{n}$, at arbitrary values of $v^{*}$. In both plots the reduced temperature is subcritical, $T/T_{c}=0.8$, thus solid lines display regions of stable phases, thin dotted horizontal lines indicate density jumps during the first-order phase transitions. The repulsion to attraction ratio $f=1.2$.
@@
Table 1: Effect of changing the reduced cell volume, $v^{*}$, on critical values of the chemical potential, $\tilde{\mu}_{c}$, and the pressure, $\tilde{P}_{c}$, corresponding to the first three phase transitions in the sequence denoted by the superscripts (1), (2), (3) respectively. The repulsion to attraction ratio $f=1.2$.
@@ 4 Impact of the interaction intensity
A closer examination of the temperature–density phase diagrams reveals a consistency in the location of the coexistence curves with respect to the reduced particle number density $\bar{n}$, regardless of the values of $f$. Specifically, the critical points occur at approximately the same density values: at the first critical point $\bar{n}\approx 0.5$, at the second $\bar{n}\approx 1.5$, and at the third $\bar{n}\approx 2.5$ (see also Table 2). Depending on the temperature, the stable phases are observed in the following density ranges: Phase I is in the reduced density range between 00 and $0.5$, Phase II between $0.5$ and $1.5$, Phase III between $1.5$ and $2.5$, and Phase IV between $2.5$ and $3.5$. In the context of the cell model, the reduced density represents the average number of particles per cell. Therefore, it is natural that the interaction intensity does not influence the statistical properties of the model.
@@ 5 Conclusions
In the present study, we establish the roles of the two key parameters used in formulating the model. The first parameter is the volume of the cubic cell, $v$, into which the total system volume is conventionally divided. We have shown that this parameter does not influence the phase transition process. In particular, the coordinates of the critical point remain unchanged, and the equation of state in pressure-temperature-density terms retains a universal form. Additionally, the steepness of the coexistence curves between the first and second phases remains unaffected.
The second parameter significantly influences the equation of state, the critical point parameters, and several other important characteristics. This parameter represents the ratio of repulsive to attractive intensities between particles, $f$, which, in our case, is a Curie-Weiss-type interaction (2). According to [26], to ensure the stability of interaction, this ratio cannot be less than one, $f\geq 1$. That is, attraction cannot exceed repulsion, otherwise, the system will tend to collapse. Our results confirm that increasing the repulsion-to-attraction ratio raises the critical temperature. However, when this ratio exceeds 3, $f\geq 3$, the critical temperature stabilizes and no longer increases. In this case, the critical temperatures of each phase transition in the sequence are equal. As $f$ increases, the pressures at which the phase transitions occur also rise, particularly for the second and third transitions. Regardless of the repulsion-to-attraction ratio, the critical points occur at approximately the same reduced densities: 0.5, 1.5, and 2.5. This consistency reflects the nature of the reduced density as the average number of particles per cell in the model, which is unaffected by changes in $f$.
In summary, varying the cell volume $v$ does not affect the phase diagrams in any way, whereas the ratio of repulsive to attractive interaction intensities between particles $f$ affects the phase diagrams quantitatively, but does not lead to qualitative changes in the overall phase behavior of the cell model.

### v2

@@ Abstract
We investigate how varying two microscopic parameters – cell volume and the ratio between repulsion and attraction intensities – affect the phase behavior of a cell model with Curie-Weiss-type interaction. The analysis is based on an exact solution previously derived for this model in the grand canonical ensemble. At sufficiently low temperatures, the cell model exhibits multiple first-order phase transitions. By varying the cell volume and the repulsion-to-attraction ratio, we represent a quantitative comparison of the chemical potential and pressure isotherms, along with the pressure-temperature and temperature-density phase diagrams. Our results demonstrate that altering these microscopic parameters induces quantitative changes in the phase diagrams of the cell model.
@@ 1 Introduction
The analysis of fluid models using simplified interaction potentials is essential for understanding the connection between microscopic interactions and macroscopic phase behavior. Fluid phase diagrams can be significantly altered by even small changes in interaction parameters, such as well depth, range, or softness of repulsion as shown by computer modeling and theoretical investigations employing statistical mechanics, integral equations, and perturbation theories [1]. Computer simulation studies demonstrate that decreasing the range of attraction in a system of hard spheres with an attractive Yukawa potential [2, 3] as well as in Lennard-Jones like models [4, 5] leads to the vanishing of stable fluid-fluid transitions and remaining stable fluid-solid transitions. Moreover, reducing the interaction range flatten out the vapor-liquid coexistence curves and shrinks the range of temperatures for a stable liquid phase in attractive hard-core Yukawa fluids, as shown by computer modeling [6, 7, 8], reference hypernetted chain integral equation theory [9], and perturbative density functional theory [10]. The same effect is also observed in square-well fluids when decreasing the range of the potential well [11], in systems with interactions modeled by the approximate non-conformal theory potentials when weakening the attractive softness [12], in Mie fluids when either increasing repulsion exponent [13, 14] or increasing dispersive exponent [14], in fluids with modified Buckingham exponential-6 potential when increasing repulsion intensity [15, 16].
@@ 3 Effect of changing the cell size
In this Section, we study the influence of changing the reduced cell volume, $v^{*}$, on the phase behavior of the cell model. In the framework of the cell model the cell volume is a constant quantity. Moreover in the thermodynamic limit the average number of particles $\langle N\rangle\rightarrow\infty$, the total volume of the system $V\rightarrow\infty$, but $v={\mathrm{const}}$. Thus our investigation means the quantitative comparison of the values that the reduced particle number density, $\bar{n}$, the chemical potential, $\tilde{\mu}$, and the pressure, $\tilde{P}$, can take, in the cases we set $v^{*}=1$, $v^{*}=2$, $v^{*}=5$, $v^{*}=10$, $v^{*}=15$.
@@
Figure 1: Part a: influence of varying the cell size, $v^{*}$, on isotherms of the chemical potential, $\tilde{\mu}$, as a function of the reduced density, $\bar{n}$: black line corresponds to $v^{*}=1$, red line – $v^{*}=2$, blue line – $v^{*}=5$, green line – $v^{*}=10$, purple line – $v^{*}=15$. Part b: isotherm of the pressure, $\tilde{P}$, as a function of the reduced density, $\bar{n}$, at arbitrary values of $v^{*}$. In both plots the reduced temperature is subcritical, $T/T_{c}=0.8$, thus solid lines display regions of stable phases, thin dotted horizontal lines indicate density jumps during the first-order phase transitions. The repulsion to attraction ratio $f=1.2$.
@@
Table 1: Effect of varying the reduced cell volume, $v^{*}$, on critical values of the chemical potential, $\tilde{\mu}_{c}$, and the pressure, $\tilde{P}_{c}$, corresponding to the first three phase transitions in the sequence denoted by the superscripts (1), (2), (3) respectively. The repulsion to attraction ratio $f=1.2$.
@@ 4 Impact of the interaction intensity
A closer examination of the temperature–density phase diagrams reveals a consistency in the location of the coexistence curves with respect to the reduced particle number density $\bar{n}$, regardless of the values of $f$. Specifically, the critical points occur at approximately the same density values: at the first critical point $\bar{n}\approx 0.5$, at the second $\bar{n}\approx 1.5$, and at the third $\bar{n}\approx 2.5$ (see also Table 2). Depending on the temperature, the stable phases are observed in the following density ranges: Phase I is in the reduced density range between 00 and $0.5$, Phase II between $0.5$ and $1.5$, Phase III between $1.5$ and $2.5$, and Phase IV between $2.5$ and $3.5$. In the context of the cell model, the reduced density represents the average number of particles per cell.
@@ 5 Conclusions
In the present study, we establish the roles of the two key parameters used in formulating the cell model. The first parameter is the volume of the cubic cell, $v$, into which the total system volume is conventionally divided. We have shown that this parameter has no influence on the phase transition characteristics. In particular, the equation of state in pressure-temperature-density terms retains a universal form, and the coordinates of the critical point remain unchanged. However, this parameter has impact on the values of the chemical potential, $\bar{\mu}$.
The second parameter significantly influences the equation of state, the critical point parameters, and several other important characteristics. This parameter represents the ratio of repulsive to attractive intensities between particles, $f$, which, in our case, is a Curie-Weiss-type interaction (2). According to [26], to ensure the stability of interaction, this ratio cannot be less than one, $f\geq 1$. That is, attraction cannot exceed repulsion, otherwise, the system will tend to collapse. In the present study we analyzed the range $1<f\leq 10$. Our results confirm that increasing the repulsion-to-attraction ratio lowers the critical temperature. However, when this ratio exceeds 3, $f>3$, the critical temperature saturates and no longer decreases. In this case, the critical temperatures of each phase transition in the sequence are equal. As $f$ increases, the pressures at which the phase transitions occur also rise, particularly for the second and third transitions. However, as $f$ increases, the critical points continue to occur at approximately the same reduced densities as they do at lower values of $f$. This consistency reflects the nature of the reduced density as the average number of particles per cell in the model, which remains almost unaffected by changes in $f$.
In summary, varying the cell volume, $v$, does not affect the pressure-temperature-density phase diagrams. In contrast, increasing the ratio of repulsive to attractive interaction intensities between particles, $f$, influences the phase diagrams quantitatively, for example, by expanding the pressure range over which stable phases exist or by decreasing the critical temperature until it saturates, resulting in equal critical temperatures for all phase transitions in the sequence. However, changes in $f$ do not lead to qualitative alterations in the phase behavior of the cell model: phenomena such as the disappearance of phase transitions or the emergence of triple points are not observed.
Acknowledgement
Yuriy O. Plevachuk is supported by the Office of Government of Slovakia, project no. 09I03-03-V01-00047.